# Beyond Multiple Choice: A Hybrid Framework for Unifying Robust Evaluation and Verifiable Reasoning Training

### v1

@@ 1 Introduction
As large language and multimodal models (Anthropic, 2025; OpenAI, 2025; Bai et al., 2025; OpenAI, 2023; Google, 2025; Chen et al., 2025; Pei et al., 2025) increasingly tackle diverse real-world tasks, the demand for reliable and scalable evaluation has grown significantly. MCQA is convenient because restricting outputs simplifies scoring (Moore et al., 2023) across language (Hendrycks et al., 2020; Wang et al., 2024a) and vision language benchmarks (Yue et al., 2023; 2024; Liu et al., 2023; Kembhavi et al., 2016; Zhang et al., 2024; Hao et al., 2025).
However, MCQA departs from real‑world usage where answers are usually open-ended (Lyu et al., 2024), while the predefined options encourage selection heuristics (Zheng et al., 2023; Balepur et al., 2024) rather than genuine understanding. To quantify the unreliability of MCQA for evaluation and verification, we conduct multiple experiments: (1) When options are added to the questions in an open-form benchmark, the accuracy metrics can be greatly boosted; (2) In MCQA benchmarks, When the ground-truth option is perturbed, or replaced with ‘None of the above’, model behavior degrades. These patterns indicate that the MCQA metrics are heavily dependent on the option set, rather than solely on the knowledge and skills required in the question stem. This fragility matters because many visual reasoning datasets used for outcome-based RFT have included large proportions of MCQA data. We find that training on MCQA increases multiple-choice accuracy metrics but hurts open-form generalization, widening the gap between the two evaluation settings. In other words, this reward encourages shortcuts tied to options rather than transferable knowledge or reasoning (See Figure 1).
@@
Based on ReVeL, we also rewrite 20k MCQA examples into OpenQA and perform GRPO‑based RFT on Qwen2.5‑VL‑3B/7B. Models trained with ReVeL‑OpenQA match MCQA accuracy on choice benchmarks while improving OpenQA accuracy by about six percentage points, demonstrating higher data efficiency and stronger robustness than MCQA‑based training. With a modest data, OpenQA‑trained 7B models also exceed the counterparts trained on open‑source data recipes such as VL‑Rethinker‑7B (Wang et al., 2025a), R1‑OneVision‑7B (Yang et al., 2025), and Mixed‑R1‑7B (Xu et al., 2025a) on open‑ended evaluation. In summary, our contributions are threefold:
@@ 2.1 Adding options to open-ended benchmarks
Setup. We start from two recent benchmarks that expect free-form answers from an LLM or VLM: SimpleQA (Wei et al., 2024) and VisualSimpleQA (Wang et al., 2025b). We convert each question into an MCQA variant (SimpleQA‑Choice / VisualSimpleQA‑Choice) by retaining the ground-truth answer and adding five plausible distractors via a human‑in‑the‑loop procedure with GPT‑4.1. This conversion preserves the original semantics, but the metrics may be affected by random guessing. Therefore, besides accuracy, we also report a random‑guessing upper bound:
@@ 2.2 Replacing GT with None-of-the-Above
Another way to test the target knowledge or reasoning skill is to replace the ground-truth option with an option to abstain: ‘None of the above (is correct)’ (NOTA), after shifting the remaining false options frontwards. We conduct such an experiment on MMLU-Pro (Wang et al., 2024a) and MMMU (Yue et al., 2023), the most popular MCQA benchmarks for LLM and VLM evaluation.
@@ 2.4 RFT on MCQA hurts open-ended QA
Finally, we study training effects by utilizing reinforcement fine-tuning on MCQA data and evaluating on both MCQA and their open counterparts described in Section 2.3. We use the popular GRPO algorithm (Shao et al., 2024) in this work for RFT experiments. RFT on MCQA improves MCQA scores but degrades open-ended performance, thereby widening the MCQA–OpenQA gap. For example, on MMMU, the gap grows for both 3B and 7B models; similar trends hold on EMMA (see Table 2). This indicates that the verifiable reward under MCQA may overfit to option-specific heuristics rather than transferable reasoning.
@@ 3.2 Benchmarks and rewriting coverage
We evaluate ReVeL on four major multimodal benchmarks, including EMMA, MMMU, MME-RealWorld and MMLU-Pro. EMMA (Hao et al., 2025) targets multimodal reasoning in STEM, emphasizing visual–textual integration; we focus on the physics and chemistry subsets for domain-specific evaluation. MMMU (Yue et al., 2023) assesses college-level, multi-discipline reasoning across six domains with diverse image types; we use its 900-question validation set. MME-RealWorld (Zhang et al., 2024) offers large-scale, high-quality, real-world tasks with greater difficulty; we adopt its “Lite” subset of 1,700 questions. MMLU-Pro (Wang et al., 2024a) is a more challenging variant of MMLU, incorporating reasoning-oriented questions, ten-choice answers, and cleaner data. We sample 1,000 questions for evaluation.
@@ 5 Related work
Multiple-choice question answering (MCQA) has been a popularly used assessment tool for ages due to simplified grading (Simkin and Kuechler, 2005; Dufresne et al., 2002; Paxton, 2000; Balepur et al., 2025; Alzahrani et al., 2024; Pezeshkpour and Hruschka, 2023). This convenience led to its wide adoption for evaluation of large language models (Hendrycks et al., 2020; Wang et al., 2024a), and in particular vision-language models (Yue et al., 2023; Clark et al., 2018; Yue et al., 2024; Liu et al., 2024) because of more diverse wording choices in describing many visual concepts or scenes. However, MCQA has many shortcuts. Performance can drop dramatically simply from changing an option’s placement(Zheng et al., 2023; Molfese et al., 2025). While mitigation strategies—such as better distractors, more options, randomized order, or ’select all that apply’ formats (Zhang et al., 2025; Yu et al., 2024; Zheng et al., 2023; Zhou et al., 2024; Xu et al., 2025b) reduced some biases. And models typically cannot reject all options when the correct answer is absent (G’oral et al., 2024; Tam et al., 2025). Some recent work has shown that reasoning models are good at exploiting the information in the options, implying the performance may be inflated (Balepur et al., 2024; Raman et al., 2025). Recognizing these issues, the community’s shift to open-ended evaluation faces its own challenges. Rule-based, short-answer benchmarks (xAI, 2024; Wang et al., 2024b) are limited in scope, while general open-ended formats rely on an LLM-as-a-judge. Furthermore, simply remove options must discard a significant portion of unsuitable items and still depend on an LLM-Judge for evaluation (Myrzakhan et al., 2024). These works analyse the flaws of MCQA but do not try to propose a method to mitigate these shortcuts. These analyses focus on identifying the flaws of MCQA rather than proposing systematic mitigation strategies.
Multimodal reinforcement learning: Many visual reasoning datasets are predominantly designed in an MCQA format. For instance, earlier datasets such as ScienceQA (Lu et al., 2022), AI2D (Kembhavi et al., 2016), Geometry3K (Lu et al., 2021), and GeoQA-Plus (Chen et al., 2021) are entirely formed by multiple-choice questions. This trend continues in recent MLLMs designed for general-purpose reasoning, such as Mixed-R1 (Xu et al., 2025a), R1-OneVision (Yang et al., 2025), and VL-Rethinker(Wang et al., 2025a), which all employ a considerable proportion of choice-based items, accounting for 43%, 80%, and 45% of their data, respectively. Our work is built on these visual reasoning datasets and explores open-form rewriting from those MCQA samples.
@@ 6 Limitations
We acknowledge several limitations in our proposed pipeline. First, the rewriting and classification phases, while highly accurate, are not perfect and may occasionally introduce errors. Hopefully such errors could diminish when the LLM components are getting stronger and stronger in the future. Second, our work focuses on converting the format of evaluation to be more robust and efficient, without addressing the inherent fallibility of the LLM-judge itself. Issues such as positional bias, verbosity bias, or factual inaccuracies within the LLM-judge (Chen et al., 2024) are orthogonal to our contribution. We deliberately sidestep some of these known issues; for instance, questions in the EMMA dataset requiring the validation of SMILES chemical structures were intentionally converted to a Per-Option Verification format. This leverages rule-based checking and avoids relying on an LLM-judge for a domain-specific task, thereby mitigating a potential failure point of LLM-based evaluation. There are several directions for future research. One key avenue is to extend our framework beyond QA to other NLP tasks, such as long-form generation, where evaluation remains a major challenge. Finally, developing adaptive evaluation systems that can dynamically choose the most appropriate and cost effective judging mechanism based on the question’s complexity and the model’s response would be a valuable next step.

### v2

# Beyond Multiple Choice: Verifiable OpenQA for Robust Vision-Language RFT

@@ 1 Introduction
As large language and multimodal models (Anthropic2024Claude4; GPT-5; Bai2025Qwen25VLTR; openai2023gpt4v; Google2025Gemini25pro; Chen2025InternVL2.5; Pei2025EgoThinker) increasingly tackle diverse real-world tasks, the demand for reliable and scalable evaluation has grown significantly. MCQA is convenient because restricting outputs simplifies scoring (Moore2023AssessingMCQ) across language (Hendrycks2020MMLU; Wang2024MMLUPro) and vision language benchmarks (Yue2023MMMU; Yue2024MMMUPro; Liu2023MMBench; Kembhavi2016AI2D; Zhang2024MMERealWorld; Hao2025EMMA).
However, MCQA departs from real‑world usage where answers are usually open-ended (Lyu2024BeyondPro), while the predefined options encourage selection heuristics (Zheng2023LMSelectBias; Balepur2024ArtifactsOA) rather than genuine understanding. To quantify the unreliability of MCQA for evaluation and verification, we conduct multiple experiments: (1) When options are added to the questions in an open-form benchmark, the accuracy metrics can be greatly boosted; (2) In MCQA benchmarks, When the ground-truth option is perturbed, or replaced with ‘None of the above’, model behavior degrades. These patterns indicate that the MCQA metrics are heavily dependent on the option set, rather than solely on the knowledge and skills required in the question stem. This fragility matters because many visual reasoning datasets used for outcome-based RFT have included large proportions of MCQA data. We find that training on MCQA increases multiple-choice accuracy metrics but hurts open-form generalization, widening the gap between the two evaluation settings. In other words, this reward encourages shortcuts tied to options rather than transferable knowledge or reasoning (See Figure 1).
@@
Based on ReVeL, we also rewrite 20k MCQA examples into OpenQA and perform GRPO‑based RFT on Qwen2.5‑VL‑3B/7B. Models trained with ReVeL‑OpenQA match MCQA accuracy on choice benchmarks while improving OpenQA accuracy by about six percentage points, demonstrating higher data efficiency and stronger robustness than MCQA‑based training. With a modest data, OpenQA‑trained 7B models also exceed the counterparts trained on open‑source data recipes such as VL‑Rethinker‑7B (Wang2025VLRethinker), R1‑OneVision‑7B (Yang2025R1Onevision), and Mixed‑R1‑7B (Xu2025MixedR1) on open‑ended evaluation. In summary, our contributions are threefold:
@@ 2.1 Adding options to open-ended benchmarks
Setup. We start from two recent benchmarks that expect free-form answers from an LLM or VLM: SimpleQA (Simpleqa) and VisualSimpleQA (Wang2025VisualSimpleQA). We convert each question into an MCQA variant (SimpleQA‑Choice / VisualSimpleQA‑Choice) by retaining the ground-truth answer and adding five plausible distractors via a human‑in‑the‑loop procedure with GPT‑4.1. This conversion preserves the original semantics, but the metrics may be affected by random guessing. Therefore, besides accuracy, we also report a random‑guessing upper bound:
@@ 2.2 Replacing GT with None-of-the-Above
Another way to test the target knowledge or reasoning skill is to replace the ground-truth option with an option to abstain: ‘None of the above (is correct)’ (NOTA), after shifting the remaining false options frontwards. We conduct such an experiment on MMLU-Pro (Wang2024MMLUPro) and MMMU (Yue2023MMMU), the most popular MCQA benchmarks for LLM and VLM evaluation.
@@ 2.4 RFT on MCQA hurts open-ended QA
Finally, we study training effects by utilizing reinforcement fine-tuning on MCQA data and evaluating on both MCQA and their open counterparts described in Section 2.3. We use the popular GRPO algorithm (Shao2024DeepSeekMath) in this work for RFT experiments. RFT on MCQA improves MCQA scores but degrades open-ended performance, thereby widening the MCQA–OpenQA gap. For example, on MMMU, the gap grows for both 3B and 7B models; similar trends hold on EMMA (see Table 2). This indicates that the verifiable reward under MCQA may overfit to option-specific heuristics rather than transferable reasoning.
@@ 3.2 Benchmarks and rewriting coverage
We evaluate ReVeL on four major multimodal benchmarks, including EMMA, MMMU, MME-RealWorld and MMLU-Pro. EMMA (Hao2025EMMA) targets multimodal reasoning in STEM, emphasizing visual–textual integration; we focus on the physics and chemistry subsets for domain-specific evaluation. MMMU (Yue2023MMMU) assesses college-level, multi-discipline reasoning across six domains with diverse image types; we use its 900-question validation set. MME-RealWorld (Zhang2024MMERealWorld) offers large-scale, high-quality, real-world tasks with greater difficulty; we adopt its “Lite” subset of 1,700 questions. MMLU-Pro (Wang2024MMLUPro) is a more challenging variant of MMLU, incorporating reasoning-oriented questions, ten-choice answers, and cleaner data. We sample 1,000 questions for evaluation.
@@ 5 Related work
Multiple-choice question answering (MCQA) has been a popularly used assessment tool for ages due to simplified grading (Simkin2005MultipleChoice; makingsenseMCQ; LinguisticMCQ; Balepur2025FlawsMCQ; alzahrani-etal-2024-MCQSensitivity; SensitivitytoTheOrder). This convenience led to its wide adoption for evaluation of large language models (Hendrycks2020MMLU; Wang2024MMLUPro), and in particular vision-language models (Yue2023MMMU; arcbench; Yue2024MMMUPro; Liu2024IIBench) because of more diverse wording choices in describing many visual concepts or scenes. However, MCQA has many shortcuts. Performance can drop dramatically simply from changing an option’s placement(Zheng2023LMSelectBias; Molfese2025inconMultiChoice). While mitigation strategies—such as better distractors, more options, randomized order, or ’select all that apply’ formats (Zhang2025AutomatedOpen2Choice; yu-etal-2024-KnowledgeEnhancingDistractor; Zheng2023LMSelectBias; zhou-etal-2024-revisiting; Xu2025SATABENCH) reduced some biases. And models typically cannot reject all options when the correct answer is absent (Goral2024IncorrectMCQ; Tam2025NoneOPfTheAbove). Some recent work has shown that reasoning models are good at exploiting the information in the options, implying the performance may be inflated (Balepur2024ArtifactsOA; Raman2025MCQExploiter). Recognizing these issues, the community’s shift to open-ended evaluation faces its own challenges. Rule-based, short-answer benchmarks (realwordqa; Wang2024CharXiv) are limited in scope, while general open-ended formats rely on an LLM-as-a-judge. Furthermore, simply remove options must discard a significant portion of unsuitable items and still depend on an LLM-Judge for evaluation (Myrzakhan2024OpenLLMLeaderboardFM). These works analyse the flaws of MCQA but do not try to propose a method to mitigate these shortcuts. These analyses focus on identifying the flaws of MCQA rather than proposing systematic mitigation strategies.
Multimodal reinforcement learning: Many visual reasoning datasets are predominantly designed in an MCQA format. For instance, earlier datasets such as ScienceQA (Lu2022ScienceQA), AI2D (Kembhavi2016AI2D), Geometry3K (Lu2021InterGPSIG), and GeoQA-Plus (Chen2021GeoQAAG) are entirely formed by multiple-choice questions. This trend continues in recent MLLMs designed for general-purpose reasoning, such as Mixed-R1 (Xu2025MixedR1), R1-OneVision (Yang2025R1Onevision), and VL-Rethinker(Wang2025VLRethinker), which all employ a considerable proportion of choice-based items, accounting for 43%, 80%, and 45% of their data, respectively. Our work is built on these visual reasoning datasets and explores open-form rewriting from those MCQA samples.
@@ 6 Limitations
We acknowledge several limitations in our proposed pipeline. First, the rewriting and classification phases, while highly accurate, are not perfect and may occasionally introduce errors. Hopefully such errors could diminish when the LLM components are getting stronger and stronger in the future. Second, our work focuses on converting the format of evaluation to be more robust and efficient, without addressing the inherent fallibility of the LLM-judge itself. Issues such as positional bias, verbosity bias, or factual inaccuracies within the LLM-judge (Chen2024MLLMasaJudge) are orthogonal to our contribution. We deliberately sidestep some of these known issues; for instance, questions in the EMMA dataset requiring the validation of SMILES chemical structures were intentionally converted to a Per-Option Verification format. This leverages rule-based checking and avoids relying on an LLM-judge for a domain-specific task, thereby mitigating a potential failure point of LLM-based evaluation. There are several directions for future research. One key avenue is to extend our framework beyond QA to other NLP tasks, such as long-form generation, where evaluation remains a major challenge. Finally, developing adaptive evaluation systems that can dynamically choose the most appropriate and cost effective judging mechanism based on the question’s complexity and the model’s response would be a valuable next step.